OpenStudy (camerondoherty):

Im 100% sure it is either A or B...

OpenStudy (camerondoherty):

Two systems of equations are shown below: System A: 2x + y = 5 -4x + 6y = -2 System B -10x + 19y = -1 -4x + 6y = -2 Which of the following statements is correct about the two systems of equations? They will have the same solutions because the first equation of System B is obtained by adding the first equation of System A to 2 times the second equation of System A. They will have the same solution because the first equation of System B is obtained by adding the first equation of System A to 3 times the second equation of System A. The value of x for System B will be –5 times the value of x for System A because the coefficient of x in the first equation of System B is –5 times the coefficient of x in the first equation of System A. The value of x for System A will be equal to the value of y for System B because the first equation of System B is obtained by adding –12 to the first equation of System A and the second equations are identical.

OpenStudy (here_to_help15):

do u want answer or help?

OpenStudy (camerondoherty):

ummm both lol

OpenStudy (here_to_help15):

lol :)

OpenStudy (here_to_help15):

ok hold on:)

OpenStudy (camerondoherty):

*waits*

OpenStudy (triciaal):

did you find the solutions so you can compare the steps taken?

OpenStudy (here_to_help15):

ok i did math for both of them and both are y=1 x=2

OpenStudy (camerondoherty):

the solutions are the same

OpenStudy (here_to_help15):

yeah

OpenStudy (camerondoherty):

i already did tht step

OpenStudy (camerondoherty):

thats why i said its either a or b

OpenStudy (here_to_help15):

so then lets see hmmmm

OpenStudy (here_to_help15):

ok im 14 and going to 9th and im answering questions besides ur questions beyond my grade level: P

OpenStudy (camerondoherty):

kool!

OpenStudy (here_to_help15):

umm i would love to say its b but let me check again

OpenStudy (nikato):

its B.

OpenStudy (here_to_help15):

yay i was right

OpenStudy (camerondoherty):

Ok ok Thanks @Nikato

OpenStudy (nikato):

if you follow the wording for B 2x+y=5 +3(-4x+6y=-2) ---------------- you get the first equation of B which is -10x+19y=-1

OpenStudy (nikato):

but if you do A, you get somethign else

OpenStudy (here_to_help15):

well i did what i could :P : ) n_n

OpenStudy (camerondoherty):

That was what i was attempting to do but i didnt get anywhere xD

OpenStudy (camerondoherty):

Thanx again

OpenStudy (nikato):

no problem

OpenStudy (triciaal):

is this still open because you need to know why the answer is B?

OpenStudy (camerondoherty):

no it was just cause i was too lazy to close it lol @triciaal